three is the number of digits you see
their sum is between sixteen and twenty three
the number is also a multiple of four the first digit is more than five and prime
the last two digits have a some of two more than ten but the one digit is greater
what is the number

Answer by Edwin McCravy(20054)


three is the number of digits you see
their sum is between sixteen and twenty three
the number is also a multiple of four
the first digit is more than five and prime.

The first digit is 7 because that's the only
digit that is more than 5 and prime.

The statement

"their sum is between sixteen and twenty three"

does not tell if it is inclusive of 16 and 23 or 
exclusive of them.

If it's "inclusive", these are the solutions.

728 = 4×182, and 7+2+8 = 17
736 = 4×184, and 7+3+6 = 16
748 = 4×187, and 7+4+8 = 19
756 = 4×189, and 7+5+6 = 18
764 = 4×191, and 7+6+4 = 17
768 = 4×192, and 7+6+8 = 21
772 = 4×193, and 7+7+2 = 16
776 = 4×194, and 7+7+6 = 20
784 = 4×196, and 7+8+4 = 19
788 = 4×197, and 7+8+8 = 23
792 = 4×198, and 7+9+2 = 18
796 = 4×199, and 7+9+6 = 22

If it's "exclusive", omit the three red solutions.

Edwin